OpenStudy (anonymous):

Bottles of concentrated acids, such as phosphoric acid, H3PO4, often list the acid’s density and its mass percent. For example, concentrated solutions of phosphoric acid have a density of 1.69 g/mL are 85.5% H3PO4. What is the molarity of a solution of concentrated phosphoric acid?

OpenStudy (anonymous):

We use the formula .855 = grams/mL To find the Amount of grams in a 85.5% solution of H3PO4 that is 1g/1mL density, you can just create a volume I will go with 1mL, So .855 = x/1mL = 0.855grams So now we need to find out how much solution would be in a 1.69g/1mL density for an 85.5% solution of H3PO4. So we convert 1mL the total volume to grams remembering that 1.00mL = 1.00g. Thus, we have: 1.69g/mL = 1g/x = 1.00g/1.69mL = 0.592mL Now find the moles of H3PO4 in 0.855grams, 0.855g/97.9952g/mol = 0.008725mol Now find Molarity using 0.592mL after converting it to liters so, 0.008725mol/0.000592L =14.74M It has been awhile since I have had to do density problems, if anyone can confirm this that would be great.